# LLM Personas as a Substitute for Field Experiments in Method Benchmarking

### v1

@@ Conclusion.
If the swap $B\leftrightarrow B^{\prime}$ were JPC on the aggregate-only method interface in the sense of Definition 2, then (since $Q_{B}=Q_{B^{\prime}}$) the factorization would force every algorithm’s aggregate transcript law to agree under $B$ and $B^{\prime}$. The above $\mathcal{A}$ contradicts this as soon as raw votes are released. Therefore leaking micro-level information (violating (AO)) can break JPC even when the aggregate kernels match. ∎
@@
The same phenomenon occurs if the benchmark releases panel identities (or stable rater IDs): even if the aggregate score distribution $Q(\cdot\mid w)$ is unchanged, the extra identifier acts as a side channel that an adaptive algorithm can condition on, producing different submission sequences and hence different aggregate transcripts. This is exactly why (AO) is a benchmark-hygiene requirement: it rules out side channels through which the method can distinguish evaluator implementations that are otherwise identical at the aggregate level.
B.2 Practical realism of algorithm-blind evaluation (AB)
@@
Assumption 2 is a protocol requirement: conditional on the submitted artifact $w$, the distribution of the returned aggregate feedback must not depend on who submitted $w$ or how it was produced. This is intentionally stronger than what one gets “by default” in many human-judgment settings, because there is extensive evidence that evaluations can shift when raters are exposed to provenance cues (e.g., explicit labels such as “AI-generated” vs. “human-written”) or other contextual metadata beyond the content being judged (bogert2022human; osborne2025me; levitt2011hawthorne; dequidt2018demand). In our framework, such provenance dependence is precisely an (AB) violation: two identical artifacts $w$ can induce different score distributions if the evaluator observes additional information correlated with the producing algorithm. [^1] If provenance cues are literally part of the submitted artifact (e.g., the text contains “as an AI language model”), then any resulting penalty is not an (AB) violation, because it is a function of $w$ itself. The problematic case for (AB) is when the evaluator is shown extra metadata (model name, submitter identity, algorithm label, timestamp-based context, etc.) that is not a function of $w$.
How realistic is (AB) in practice?
@@
(AB) is best viewed as a design target that is often approximately achievable, but not automatic. It is most realistic in settings where the outcome is behavioral and passively recorded (classic online A/B tests), since users are typically not told which algorithm produced what they see and outcomes like click-through or conversion are not direct subjective judgments (kohavi2020trustworthy). By contrast, in explicit rating / preference-judgment pipelines (crowd or expert), (AB) is fragile: even minimal provenance cues (labels, branding, prior beliefs about model quality, expectations about “humans vs. AI”) can create systematic shifts in ratings for the same artifact, violating (AB). This fragility is also consistent with the motivation for anonymous pairwise evaluation interfaces used in large-scale leaderboards (chiang2024chatbotarena) and with documented ordering/labeling effects in LLM-judge-style protocols (dominguezolmedo2024questioning).
@@
The key takeaway is that (AB) should not be read as an empirical claim about human invariance; it is a benchmark hygiene constraint that must be enforced (or at least audited) by the benchmark implementation.
Practical implementation choices.
Below are practical implementation choices that make (AB) substantially more plausible in both human and persona/LLM-judge benchmarks.
(1) Strict provenance anonymization (double-blind by default).
Ensure that evaluators never see submitter- or algorithm-identifying metadata:
  • Strip model names, system prompts, training/provenance tags, organization branding, and any UI elements that reveal variant identity.
  • Use random opaque identifiers (e.g., “Response X” and “Response Y”) and keep the mapping to algorithms on the server side only.
  • Avoid showing historical performance, leaderboards, or any signals that could prime raters about quality.
This is the direct operational interpretation of (AB): hold fixed everything visible to the evaluator except $w$.
(2) Canonicalize artifacts to remove accidental side channels.
Even if you do not explicitly reveal provenance, formatting can leak it. To reduce accidental algorithm-identifying cues unrelated to substantive quality:
  • Normalize formatting (whitespace, markdown rendering, typography, citation style, boilerplate).
  • Remove platform-inserted headers/footers (timestamps, device type, language tags) unless they are intended to be part of $w$.
  • If appropriate for the task, enforce a standardized output schema (length limits, structured fields), so that “style tells” are less tied to a specific submitter identity.
(3) Randomized and counterbalanced presentation.
Ordering effects can induce history dependence that behaves like a provenance effect at the interface:
  • Randomize left/right assignment in pairwise comparisons, and counterbalance so each artifact appears equally often in each position (dominguezolmedo2024questioning).
  • Randomize the sequence of items shown to an evaluator; avoid blocks that group outputs by algorithm.
  • If using repeated judgments, randomize task assignment so the evaluator cannot infer that two items came from the same submitter.
(4) Evaluation pipeline versioning and stability.
A common (AB) failure mode is a slowly drifting evaluator: rubrics change, judge models are updated, or panel composition shifts in a way correlated with submission time.
  • Freeze the rubric, instructions, and (for LLM judges) the model version and decoding parameters during a benchmark run.
  • Treat any change as a new benchmark version, and report results separately across versions.
This helps ensure that $Q(\cdot\mid w)$ is time-invariant within the horizon where algorithms adapt.
(5) Caching and seeding policies that do not create algorithm-dependent behavior.
Caching is useful (cost, latency, reproducibility) but can silently break (AB) if implemented in a history- or submitter-dependent way. Two safe patterns are:
  • Fresh-evaluation mode (i.i.d. draws): do not cache micro-responses; each call redraws the panel and re-elicits judgments. This best matches the reduced-form model in (5) and supports repeated sampling for variance estimation (Section 5).
  • Deterministic-evaluation mode (hash-seeded): if you must cache, cache by artifact only (e.g., key by a cryptographic hash of a canonicalized $w$) and derive all benchmark randomness (panel sampling and LLM sampling seeds) deterministically from that hash. Then the returned score is a function of $w$ alone, which is still compatible with (AB), but repeated submissions of the same $w$ no longer provide independent samples.
Crucially, avoid caching keyed by submitter identity, rate limits that depend on algorithm family, or batching rules that treat some submitters differently.
(6) Auditing (AB) directly via placebo provenance tests.
Because provenance effects are empirically documented, it is good practice to measure residual (AB) violations:
  • Show the same artifact under different labels (or different UI wrappers) and estimate the induced score shift. Any detectable shift is direct evidence against exact (AB).
  • Repeat the same evaluation at different times / panel slices to detect drift-induced dependence.
These audits can be reported as a benchmark diagnostic alongside main results.
B.3 End-to-end example: prompt-edit distance, $\widehat{\kappa}_{Q}$, and predicted sample size
We instantiate $(d_{\mathcal{W}},r)$ for prompt-edit-based instruction tuning, where the artifact $w\in\mathcal{W}$ is a system prompt. We restrict attention to a neighborhood $\mathcal{W}_{0}=\{w:\ d_{\mathcal{W}}(w,w_{\mathrm{base}})\leq R\}$ around a baseline prompt.
Prompt-edit metric $d_{\mathcal{W}}$.
We canonicalize each prompt into an ordered list of instruction clauses (bullets / numbered constraints). We then define $d_{\mathcal{W}}(w,w^{\prime})$ as Levenshtein distance between clause sequences with unit-cost insert/delete/substitute operations. This matches the atomic edits developers make when iterating on prompts.
Resolution $r$.
A natural minimal meaningful iteration for this family is one atomic edit, so we set $r=1$ (i.e., we require the benchmark to resolve prompts that differ by at least one edit).
Estimating $\kappa_{Q}$ from pilot runs.
We sample $M$ candidate prompts in $\mathcal{W}_{0}$ by applying random valid edits. For each prompt $w^{(i)}$ we run $n_{0}$ independent evaluation calls to obtain aggregate scores $o^{(i)}_{1},\dots,o^{(i)}_{n_{0}}\sim Q_{\mathrm{pers}}(\cdot\mid w^{(i)})$. We estimate $\widehat{\mu}_{i}$ and $\widehat{\sigma}_{i}^{2}$ and compute $\widehat{\mathrm{SNR}}_{ij}=(\widehat{\mu}_{i}-\widehat{\mu}_{j})^{2}/(\widehat{\sigma}_{i}^{2}+\widehat{\sigma}_{j}^{2})$ for all pairs with $d_{\mathcal{W}}(w^{(i)},w^{(j)})\geq r$. We take a conservative lower bound $\widehat{\kappa}_{Q}^{\mathrm{LCB}}$ as the $q$-quantile (e.g. $q=0.05$) of these SNR values.
Predicted sample size and validation.
Lemma 5.2 predicts that $n_{\mathrm{req}}=\lceil\frac{2}{\widehat{\kappa}_{Q}^{\mathrm{LCB}}}\log(1/\delta)\rceil$ independent evaluation calls suffice to misrank a $d_{\mathcal{W}}\geq r$ improvement with probability at most $\delta$. We can validate by selecting hard prompt pairs near the estimated worst-case SNR, running many repeated comparisons for varying $n$, and verifying that the empirical misranking rate crosses below $\delta$ near $n_{\mathrm{req}}$ and decays approximately as $\exp(-n\widehat{\kappa}_{Q}^{\mathrm{LCB}}/2)$.
B.4 Extension to heteroscedastic Gaussian reduced-form kernels

### v2

@@ Conclusion.
If the swap $B\leftrightarrow B^{\prime}$ were JPC on the aggregate-only method interface in the sense of Definition 2, then (since $Q_{B}=Q_{B^{\prime}}$) the factorization would force every method’s aggregate transcript law to agree under $B$ and $B^{\prime}$. The above $\mathcal{A}$ contradicts this as soon as raw votes are released. Therefore leaking micro-level information (violating (AO)) can break JPC even when the aggregate kernels match. ∎
@@
The same phenomenon occurs if the benchmark releases panel identities (or stable rater IDs): even if the aggregate score distribution $Q(\cdot\mid w)$ is unchanged, the extra identifier acts as a side channel that an adaptive method can condition on, producing different submission sequences and hence different aggregate transcripts. This is exactly why (AO) is a benchmark-hygiene requirement: it rules out side channels through which the method can distinguish evaluator implementations that are otherwise identical at the aggregate level.
B.2 Practical realism of method-blind evaluation (MB)
@@
Assumption 2 is a protocol requirement: conditional on the submitted artifact $w$, the distribution of the returned aggregate feedback must not depend on who submitted $w$ or how it was produced. This is intentionally stronger than what one gets “by default” in many human-judgment settings, because there is extensive evidence that evaluations can shift when raters are exposed to provenance cues (e.g., explicit labels such as “AI-generated” vs. “human-written”) or other contextual metadata beyond the content being judged (Bogert et al., 2022; Osborne and Bailey, 2025; Levitt and List, 2011; de Quidt et al., 2018). In our framework, such provenance dependence is precisely an (MB) violation: two identical artifacts $w$ can induce different score distributions if the evaluator observes additional information correlated with the producing method. [^1] If provenance cues are literally part of the submitted artifact (e.g., the text contains “as an AI language model”), then any resulting penalty is not an (MB) violation, because it is a function of $w$ itself. The problematic case for (MB) is when the evaluator is shown extra metadata (model name, submitter identity, method label, timestamp-based context, etc.) that is not a function of $w$.
How realistic is (MB) in practice?
@@
(MB) is best viewed as a design target that is often approximately achievable, but not automatic. It is most realistic in settings where the outcome is behavioral and passively recorded (classic online A/B tests), since users are typically not told which method produced what they see and outcomes like click-through or conversion are not direct subjective judgments (Kohavi et al., 2020). By contrast, in explicit rating / preference-judgment pipelines (crowd or expert), (MB) is fragile: even minimal provenance cues (labels, branding, prior beliefs about model quality, expectations about “humans vs. AI”) can create systematic shifts in ratings for the same artifact, violating (MB). This fragility is also consistent with the motivation for anonymous pairwise evaluation interfaces used in large-scale leaderboards (Chiang et al., 2024) and with documented ordering/labeling effects in LLM-judge-style protocols (Dominguez-Olmedo et al., 2024).
@@
The key takeaway is that (MB) should not be read as an empirical claim about human invariance; it is a benchmark hygiene constraint that must be enforced (or at least audited) by the benchmark implementation.
B.3 Extension to heteroscedastic Gaussian reduced-form kernels
@@
Appendix C Experiment: Discriminability Calibration for the Persona-Based Ad Benchmark
This experiment provides a proof-of-concept discriminability audit for persona-simulation evaluation environment used in demonstrating the effectiveness of TextBO (Kang and Yoganarasimhan, 2025), a prompt optimization based self-improving AI method. Specifically, we focus on the ad optimization experimentation proposed in this paper.
In the ad optimization experiment, TextBO iteratively improves the prompt for ad image generation that is fed to an image model to produce the ad creative, using evaluation feedback to decide which prompt edits to keep. Concretely, it runs an automated prompt improvement loop: write a prompt $\rightarrow$ generate an ad image $\rightarrow$ evaluate it on a target audience $\rightarrow$ reflection $\rightarrow$ edit the prompt and repeat.
As in Kang and Yoganarasimhan (2025), we consider eight synthetic ad campaign scenarios that cover a diverse range of products across distinct categories, each defined by a creative brief that outlines the strategic and creative direction (see Web Appendix C.1 of Kang and Yoganarasimhan (2025) for the full creative briefs):
  • Scenario 1: “GreenBite,” a new plant-based burger patty.
  • Scenario 2: “AuraSonics X1,” high-end, noise-canceling wireless earbuds.
  • Scenario 3: “Odyssey E-SUV,” a new all-electric family SUV.
  • Scenario 4: “Oasis Eco-Lodge,” a secluded, luxury resort with beautiful natural surroundings.
  • Scenario 5: “Momentum,” a mobile-first banking app for freelancers and the gig economy.
  • Scenario 6: “MindGarden,” a subscription-based meditation and mindfulness app.
  • Scenario 7: “Aeterno,” a classic, automatic Swiss-made wristwatch with a heritage design.
  • Scenario 8: “SyncFlow,” a project management and collaboration software platform for remote teams.
For evaluation, TextBO utilizes Twin-2k-500 persona dataset (Toubia et al., 2025). For a fixed scenario and prompt, TextBO evaluates an ad image by: (i) sampling 200 personas from the Twin-2k-500 training split, (ii) conditioning a multimodal LLM judge on each persona’s survey answers, requesting an effectiveness rating on the 1–5 scale, (iii) converting the judge’s log-probabilities over {1,2,3,4,5} into a real-valued expected score per persona, and (iv) averaging over the 200 personas to produce a single scalar score. This single scalar is the only feedback used by the optimization method. We use Gemini 2.5 Flash with the same meta-prompt used in Kang and Yoganarasimhan (2025), which is provided in Figure 1.
Persona prompt for simulating ad effectiveness. SYSTEM: $\{$ You are an AI assistant. Your task is to answer the TASK as if you are the individual described in the ‘Persona Profile’ (which contains their past survey responses). Remain consistent with the persona’s past survey responses and stated characteristics. Carefully follow any instructions provided for the new question, including formatting requirements. } PERSONA DATA: $\{$
Which part of the United States do you currently live in?
Question Type: Single Choice
Options:
1 - Northeast (PA, NY, NJ, RI, CT, MA, VT, NH, ME)
2 - Midwest (ND, SD, NE, KS, MN, IA, MO, WI, IL, MI, IN, OH)
3 - South (TX, OK, AR, LA, KY, TN, MS, AL, WV, DC, MD, DE, VA, NC, SC, GA, FL)
4 - West (WA, OR, ID, MT, WY, CA, NV, UT, CO, AZ, NM)
5 - Pacific (HI, AK)
Answer: 2 - Midwest (ND, SD, NE, KS, MN, IA, MO, WI, IL, MI, IN, OH)
What is the highest level of schooling or degree that you have completed?
Question Type: Single Choice
Options:
1 - Less than high school
2 - High school graduate
3 - Some college, no degree
4 - Associate’s degree
5 - College graduate/some postgrad
6 - Postgraduate
Answer: 3 - Some college, no degree
$\ldots$ (Many other survey questions and answers) $\ldots$
Suppose you were given \$5 and had to offer to another (anonymous) person a way to split the money. The other person can either accept or reject your offer. If the other person accepts your offer, you would each receive the amount you proposed. If the other person rejects your offer, you would both receive \$0. How much would you offer to the other person?
Question Type: Single Choice
Options:
1 - \$0
2 - \$1
3 - \$2
4 - \$3
5 - \$4
6 - \$5
Answer: 3 - \$2
$\ldots$ (Many other survey questions and answers) $\ldots$
AD IMAGE: [image] TASK: Return only one item from [”1”,”2”,”3”,”4”,”5”] for ad effectiveness.
Effective Score Scale Definition:
1: Extremely Unlikely. The persona would actively ignore or be annoyed by this ad.
2: Unlikely. The persona would likely scroll past without a second thought.
3: Mediocre. It is hard to decide whether the personal would click or don’t click.
4: Likely. The persona is intrigued and has a good chance of clicking to learn more.
5: Extremely Likely. The persona is the ideal target; a click is almost certain.
No explanation. Just the score.
Figure 1: Meta-prompt for simulating the effectiveness of a given ad-persona combination.
Here, in this paper’s language, an artifact $w$ corresponds to [scenario + prompt + its generated ad image]. We denote an evaluation operation of an artifact $w$, which returns a single scalar $o\in\mathbb{R}$, as Eval(w). Repeating Eval(w) with fresh persona samples yields i.i.d. draws $o\sim Q_{\mathrm{pers}}(\cdot\mid w)$. As TextBO iteratively improves the prompt for ad image generation, we would like to answer the following question:
How many personas Eval(w) need to be confident that the improved prompt is better than the original prompt?
In this problem, $d_{W}(w,w^{\prime})$ naturally corresponds to the number of clause-level instruction edits needed to transform prompt $\pi$ into $\pi^{\prime}$. Therefore, we set the resolution $r=1$, i.e., we require the benchmark to reliably distinguish prompts that differ by at least one step of prompt improvement.
For each artifact $w\in W_{0}$, we run $n_{0}$ independent repeats of TextBO’s evaluation call:
$o_{1}(w),\dots,o_{n_{0}}(w)\;\;\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\;\;Q_{\mathrm{pers}}(\cdot\mid w),$
where each repeat resamples a fresh panel of $n_{0}=200$ personas (all other evaluation components are held fixed, including the judge prompt and model version). We compute the empirical mean and variance:
$\widehat{\mu}(w)=\frac{1}{n_{0}}\sum_{t=1}^{n_{0}}o_{t}(w),\qquad\widehat{\sigma}^{2}(w)=\frac{1}{n_{0}-1}\sum_{t=1}^{n_{0}}(o_{t}(w)-\widehat{\mu}(w))^{2}.$
For each scenario, for all pairs $w,w^{\prime}\in W_{0}$ with $d_{W}(w,w^{\prime})\geq r$, we compute the pairwise signal-to-noise proxy:
$\widehat{\mathrm{SNR}}(w,w^{\prime})\;=\;\frac{\big(\widehat{\mu}(w)-\widehat{\mu}(w^{\prime})\big)^{2}}{\widehat{\sigma}^{2}(w)+\widehat{\sigma}^{2}(w^{\prime})}.$
There are many possible combinations of $w,w^{\prime}$. Kang and Yoganarasimhan (2025) reported results for ten steps of prompt improvement from the initial prompt; we tested ten improvement variants for each step. This totals 100 TextBO prompt improvement cases per scenario. We then define a conservative discriminability estimate as a low quantile of these values, instead of taking a literal infimum:
$\widehat{\kappa}_{Q}\;=\;\mathrm{Quantile}_{q}\Big(\{\widehat{\mathrm{SNR}}(w,w^{\prime}):\;d_{W}(w,w^{\prime})\geq r\}\Big),$
with $q=0.05$. We report $\widehat{\kappa}_{Q}$ per scenario and use the minimum across scenarios as the overall conservative estimate. The predicted number of independent evaluation calls per prompt required to decide between two $r$-separated artifacts is
$n_{\mathrm{req}}\;=\;\left\lceil\frac{2}{\widehat{\kappa}_{Q}}\log\frac{1}{\delta}\right\rceil.$
Table 1 provides $\widehat{\kappa}_{Q}$ and $n_{\mathrm{req}}$ for each scenario. Across the scenarios, $n_{\text{req}}$ varies from 20 to 513, with an average of 204.38. This shows that (Kang and Yoganarasimhan, 2025)’s choice of random sampling 200 personas for training was a legitimate choice.
Table 1: $\widehat{\kappa}_{Q}$ and Sample Size for $\delta=0.05$
Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 Scenario 6 Scenario 7 Scenario 8 Average
$\widehat{\kappa}_{Q}$ 0.00508 0.02320 0.13178 0.00942 0.02955 0.13069 0.0046 0.07548 0.051225
$n_{\text{req}}$ 513 113 20 280 89 20 565 35 204.375